Fermi National Laboratory
 

Accelerator Update for July 13, 2007 - July 16, 2007

 

Friday July 13

The day shift began with D-Zero (D0) in access, with the Collider Detector at Fermilab (CDF) waiting for colliding beam, with Operations monitoring the Recycle stash (288.3E10), with the Antiproton Source (Pbar) stacking (14.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Meson MTest experiment off due to a problem with a power supply transformer failure (S:HP3US).

D0 came out of access at 8:26 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 9:02 AM.

Operations put the TeV into shot setup at 9:32 AM.

Operations established store 5551 at 11:34 AM with an initial luminosity of 207.84E30.

The Neutron Therapy Facility (NTF) reported at 12:34 PM that they had finished treating their last patient for the day.

A MI expert began a study period at 1:25 PM.

At 1:39 PM, the NuMI horn tripped off. It reset.

The NuMI experiment reported at 1:50 PM that they would be running without their horn this weekend.

Power supply experts reported at 2:13 PM that the dipole magnet (HP3US) had been successfully hipotted. However, there was still a problem with its transformer.

At 3:53 PM, NuMI tripped off due to beam in the kicker gap. It reset.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations reported at 6:08 PM that Pbar experts conducted a study during an extended Recycler transfer.

At 8:34 PM, a Pbar Accumulator RF station (ARF1) tripped off. It reset.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Saturday July 14

The midnight shift began with Operations monitoring store 5551 (62.13E30) and the stash (197.7E10), with Pbar stacking (26.4E10), and with NuMI and MiniBooNE taking beam, but with MTest still off due to a transformer problem.

At 1:56 AM, a Booster RF station (BRF2) tripped off for the first of five times in 37 minutes. It reset each time. After the last trip, operators reduced the station’s voltage.

At 5:05 AM, LRF2 tripped off. It reset.

Operations reported at 7:54 AM that the Recycler had stashed three times during the midnight shift.

A mechanical expert reported that the failed HP3US transformer was removed at 8:05 AM. They would be moving the replacement transformer in shortly.

At 9:06 AM, the MI/MiniBooNE beam permit tripped off due to a bad beam position at the MI-8 collimators for the first of two times in 22 minutes. It reset both times.

At 9:47 AM, the MI Power Supply interlock Loop tripped off with several inhibits. Operators reset most of them, but the MI-20L Perm Loop would not clear and the VCB was closed and would not open. Operators investigated.

Operations reported at 10:56 AM that experts working on HP3US transformer change investigated the MI-20L problem and replaced its PLC power supply.

Operations terminated store 5551 at 12:02 PM.

Operations put the TeV into shot setup at 12:19 PM.

At 1:57 PM, Operations reported that during shot setup the TeV lost a large portion of the protons during the helix opening. A TeV expert was called in to investigate.

Operations reported at 2:16 PM that work on the HP3US transformer was done for the day.

Operations established store 5552 at 2:23 PM with an initial luminosity of 147.44E30.

CDF and D0 reported at 2:34 PM that they were suffering form bad beam positions.

Operations reported at 8:06 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5552 was at 39.11E30, the stack was at 12.2E10, and the stash was at 120.3E10.

Operations terminated store 5552 at 10:35 PM.

Operations put the TeV into shot setup at 10:48 PM.

Sunday July 15

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:37 AM, the Linac buncher tripped off. It reset.

Operations established store 5553 at 12:56 AM with an initial luminosity of 134.41E30.

At 2:18 AM, a Linac Klystron RF station (KRF2) tripped off for the first of two times in about four hours. It reset.

Operations reported at 7:53 AM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5553 was at 59.14E30, the stack was at 37.1E10, and the stash was at 104.9E10.

Operations reported at 9:28 AM, that they had successfully powered HP3US. Operations began searching and securing the MTest enclosures.

At 9:45 AM, LRF3 tripped off on a vacuum inhibit. It would not reset. Operators called in an expert. He managed to get the station back up and running by 12:32 PM. Operations resumed sending beam to Pbar, NuMI, and MiniBooNE.

Operations established beam to MTest at 12:41 PM.

At 1:58 PM, cryo system experts at CHL reported a failing cold compressor. They brought on a backup compressor at 2:42 PM.

At 3:26 PM, Operations reported that store 5553 aborted due to a loss of communication with TeV sector C2. Operators investigated and found that a CAMAC crate in C2 was not responding due to a failed power supply. Operators had it replaced by 4:21 PM.

Operations disabled MTest for the night at 5:39 PM.

Operations put the TeV into shot setup at 5:41 PM.

Operations established store 5554 at 7:35 PM with an initial luminosity of 162.51E30.

Operations reported at 8:05 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: store 5554 was at 142.67E30, the stack was at 48.2E10, and the stash was at 7.4E10.

At 8:42 PM, NuMI tripped off due to a bad pulse of beam. It reset.

Monday July 16

The midnight shift began with Operations monitoring store 5554 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 6:18 AM, BRF2 tripped off. It reset.

Operations established beam to MTest at 6:25 AM.

Operations reported at 7:52 AM that the Recycler had stashed five times during the last 12 hours.

At 8:04 AM, LRF3 tripped off.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/16/2007    email Fermilab